# PiCO: Peer Review in LLMs based on the Consistency Optimization

### v2

@@ Appendix B Detailed Prompt for Reviewers
Appendix C Scoring Methodology
@@
Appendix D Overall Algorithm of Peer Review
The overall algorithm, as delineated in Algorithm 1, encapsulates the comprehensive process outlined in Section 2.2. This sequence commences with "Data Collection and LLMs Candidate Pool Construction," progresses through "Answer-Ranking Data Construction Based on Peer Review," advances to "Consistency Optimization," and culminates with the "Unsupervised Elimination Mechanism."
@@ Appendix D Overall Algorithm of Peer Review
Appendix E Complete Experimental Results
In Section 3.4, we both employ elimination mechanisms to cull the weakest LLMs from the ’reviewer queue’ during the evaluation process. In Figures 7 and 8, we present the results for the PEN and LIS metrics, where lower PEN scores indicate better performance, and higher LIS scores denote superior performance. It is evident that both the ’PiCO’ and PRE approaches demonstrate enhanced performance as the number of eliminated ’reviewers’ increases. In most cases, the proposed ’PiCO’ method outperforms PRE.
Figure 7: Performance comparison of the PiCO (Ours) and PRE[17] methods on the MT-Bench, Chatbot Arena, and AlpacaEval datasets, with the number of eliminated reviewers on the x-axis. The y-axis is PEN, where lower values indicate better performance.
Figure 8: Performance comparison of the PiCO (Ours) and PRE[17] methods on the MT-Bench, Chatbot Arena, and AlpacaEval datasets, with the number of eliminated reviewers on the x-axis. The y-axis is LIS, where upper values indicate better performance.
In Section 3.5, we validate the effectiveness of the consistency assumption and compare it with the Average Performance of the Reviewer Queue, i.e., employing a single LLM as the ’reviewer’ to evaluate all response pairs and then calculating the average results of all LLMs. The comprehensive results compared with the Reviewer Queue are illustrated in Table3, Figure 9, 10 and 11, revealing that in the full Reviewer Queue, the performance of the vast majority of LLMs is very poor, indicating that the evaluations from most LLMs are noise. However, our ’PiCO’ approach nearly matches the evaluative prowess of the pool’s most capable LLM, GPT-3.5. Remarkably, given its unsupervised nature, the ’PiCO’ method demonstrates the capability to mitigate the influence of noise, reaching the evaluation upper bound (the strongest LLM) within any given unknown LLM pool $M$, even in the absence of prior ranking information.
Table 3: Comparison of performance across three datasets using Unsupervised methods versus using single models in reviewer queue..
@@ Appendix E Complete Experimental Results
Appendix F Selected Models and Optimized Ranking
@@
We have comprehensively detailed the ranking outcomes across three distinct datasets for our comparative analysis, incorporating the optimized model rankings, names, and their respective scores. As delineated in Appendix C, the PiCO (Ours) is capable of employing various scoring mechanisms, thereby facilitating the presentation of ranking outcomes on three datasets utilizing both the Elo and Rank mechanisms. Furthermore, we have also enumerated the ranking results for PRD and PRE methodologies across the three datasets, offering a holistic view of the competitive landscape.
F.1 PiCO

### v3

@@ Appendix C Detailed Prompt for Reviewers
Appendix D Scoring Methodology
@@
Appendix E Overall Algorithm of Peer Review
The overall algorithm, as delineated in Algorithm 1, encapsulates the comprehensive process outlined in Section 2.2. This sequence commences with "Data Collection and LLMs Pool Construction," progresses through "Answer-Ranking Data Construction Based on Peer Review," advances to "Consistency Optimization," and culminates with the "Unsupervised Elimination Mechanism."
@@ Appendix E Overall Algorithm of Peer Review
Appendix F Complete Experimental Results
In Section 3.4, we both employ elimination mechanisms to cull the weakest LLMs from the ’reviewer queue’ during the evaluation process. In Figures 8 and 9, we present the results for the PEN and LIS metrics, where lower PEN scores indicate better performance, and higher LIS scores denote superior performance. It is evident that both the ’PiCO’ and PRE approaches demonstrate enhanced performance as the number of eliminated ’reviewers’ increases. In most cases, the proposed ’PiCO’ method outperforms PRE.
Figure 8: Performance comparison of the PiCO (Ours) and PRE (Chu et al., 2024) methods on the MT-Bench, Chatbot Arena, and AlpacaEval datasets, with the number of eliminated reviewers on the x-axis. The y-axis is CIN, where lower values indicate better performance.
Figure 9: Performance comparison of the PiCO (Ours) and PRE (Chu et al., 2024) methods on the MT-Bench, Chatbot Arena, and AlpacaEval datasets, with the number of eliminated reviewers on the x-axis. The y-axis is LIS, where upper values indicate better performance.
In Section 3.5, we validate the effectiveness of the consistency assumption and compare it with the Average Performance of the Reviewer Queue, i.e., employing a single LLM as the ’reviewer’ to evaluate all response pairs and then calculating the average results of all LLMs. The comprehensive results compared with the Reviewer Queue are illustrated in Table5, Figure 10, 11 and 12, revealing that in the full Reviewer Queue, the performance of the vast majority of LLMs is very poor, indicating that the evaluations from most LLMs are noise. However, our ’PiCO’ approach nearly matches the evaluative prowess of the pool’s most capable LLM, GPT-3.5. Remarkably, given its unsupervised nature, the ’PiCO’ method demonstrates the capability to mitigate the influence of noise, reaching the evaluation upper bound (the strongest LLM) within any given unknown LLM pool $M$, even in the absence of prior ranking information.
Table 5: Comparison of performance across three datasets using Unsupervised methods versus using single models in reviewer queue.
@@ Appendix F Complete Experimental Results
Appendix G Selected Models and Optimized Ranking
@@
We have comprehensively detailed the ranking outcomes across three distinct datasets for our comparative analysis, incorporating the optimized model rankings, names, and their respective scores. As delineated in Appendix D, the PiCO (Ours) is capable of employing various scoring mechanisms, thereby facilitating the presentation of ranking outcomes on three datasets utilizing both the Elo and Rank mechanisms. Furthermore, we have also enumerated the ranking results for PRD and PRE methodologies across the three datasets, offering a holistic view of the competitive landscape.
G.1 PiCO